# Scale dependence improvement of the quartic scalar field thermal effective potential in the optimized perturbation theory

### v1

@@ I Introduction
In the context of OPT, a new scheme called Renormalization Group Optimized Perturbation Theory (RGOPT) was developed to address the scale dependence problem [41, 42, 43, 44, 45, 46, 47, 48]. This method modifies the original OPT implementation and introduces a new (subtraction) term into the free energy, ensuring invariance under the renormalization group. The traditional OPT mass term, $(1-\delta)\eta$, is written as $(1-\delta)^{a}\eta$ where $a$ turns out to be a function of the RG $\beta$ and $\gamma_{m}$ coefficients $b_{0}$ and $\gamma_{0}$. These adaptations have allowed this approximation to achieve great success in investigating the symmetric phase of several models, including QCD, as mentioned earlier. However, so far the method has not yet been used, for example, in the description of phase transitions. The approach that we propose in the present paper differs fundamentally from the RGOPT by preserving the original prescription of the OPT method and improving over the scale dependence displayed by the thermodynamic quantities by applying directly the OPT resummation over the RGI improved effective potential. We termed this alternative approach the Variational Renormalization Group (VRG). We will see that by preserving the original OPT method, augmented with RG invariance properties, we can study the system’s thermodynamic properties in both the symmetric and broken phases in a systematic way.
This work is organized as follows. In Sec. II, we briefly review the application of the OPT method to our fiducial $\lambda\phi^{4}$ scalar field theory. The effective potential at finite temperature up to second order in the OPT is explicitly written down. In Sec. III, we review how the RGI method allows an improvement of the effective potential with respect to scale variations. In Sec. IV, we show how the RGI method can be modified to carry out an improvement in the context of the OPT finite-temperature effective potential, ensuring that the improvement remains consistent with the renormalization group. In Sec. V, we present our main numerical results for the thermodynamics of the quartic scalar field when using the VRG. Finally, in Sec. VI, we draw our main conclusions and perspectives for future work. Four appendices are also included where the relevant definitions and expressions are given and the relevant technical details used in our analysis are also explained.
@@ II The effective potential of the $\lambda\phi^{4}$ model in the OPT at next-to-leading order
$\mathcal{L}$ $=$ $\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi)-\frac{m^{2}}{2}\phi^{2}-\frac{\lambda}{4!}\phi^{4}-\Lambda,$ (1)
where $\Lambda$ is a (field-independent) “cosmological constant” that is necessary to deal with (zero point) vacuum energy terms [49]. The OPT prescription starts by implementing in the Lagrangian density the following replacements [29, 30, 31, 32]
@@
$\mathcal{L}^{\delta}$ $=$ $\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi)-\frac{\Omega^{2}}{2}\phi^{2}+\frac{\delta\eta^{2}}{2}\phi^{2}-\frac{\delta\lambda}{4!}\phi^{4}-\Lambda,$
@@
In practice, when applying the OPT, one can use the results obtained by standard perturbation theory up to some order $\lambda^{k}$ in the coupling and then apply the changes given by Eqs. (2) and (3), while re-expanding the result up to the desired order $\delta^{k}$. In the present application, we are mainly concerned with the finite temperature effective potential $V_{\rm eff}$ for the scalar quartic model at order-$\delta^{2}$. Applying Eqs. (2) and (3) to the perturbative ${\cal O}(\lambda^{2})$ standard result and re-expanding yields the corresponding OPT expression up to order-$\delta^{2}$ (NLO), whose contributions are represented by the Feynman diagrams shown in Fig. 1.
@@
$V_{\rm OPT}^{\delta^{2}}$ $=$ $\Lambda+\frac{1}{2}\left[m^{2}+(1-\delta)\eta^{2}\right]\varphi^{2}+\delta\frac{\lambda}{4!}\varphi^{4}-\hbar\left[\frac{\Omega^{4}\left(2L_{\Omega}+3\right)}{8\left(4\pi\right)^{2}}+\frac{J_{0,\Omega}T^{4}}{2\left(4\pi\right)^{2}}\right]$ (6)
$+$ $\delta\left\{\frac{\hbar^{2}\lambda}{8\left(4\pi\right)^{4}}\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]^{2}+\frac{\hbar\eta^{2}}{2\left(4\pi\right)^{2}}\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]\right$.
- $\left.\frac{\varphi^{2}}{2}\frac{\hbar\lambda}{2(4\pi)^{2}}\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]\right\}+\mathcal{F}_{\delta}$
$+$ $\delta^{2}\left\{-\frac{\hbar\eta^{4}}{4\left(4\pi\right)^{2}}\left[L_{\Omega}+J_{2,\Omega}\right]-\frac{\hbar^{2}\lambda}{4\left(4\pi\right)^{4}}\eta^{2}\left[L_{\Omega}+J_{2,\Omega}\right]\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]\right$.
- $\frac{\hbar^{3}\lambda^{2}}{48\left(4\pi\right)^{6}}\left[\left(\left(3L_{\Omega}+4\right)J_{1,\Omega}^{2}+J_{1,\Omega}^{2}J_{2,\Omega}+2K_{2,\Omega}+\frac{4}{3}K_{3,\Omega}\right)3T^{4}\right$.
@@
$+$ $\frac{\varphi^{2}}{2}\left[-\frac{\hbar^{2}\lambda^{2}\Omega^{2}}{4(4\pi)^{4}}\left((L_{\Omega}+1)^{2}+1+\frac{\pi^{2}}{6}\right)+\frac{\hbar\lambda\eta^{2}}{2(4\pi)^{2}}\left(L_{\Omega}+J_{2,\Omega}\right)\right$.
$+$ $\frac{\hbar^{2}\lambda^{2}}{4(4\pi)^{4}}\left(\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right)\left(L_{\Omega}+J_{2,\Omega}\right)$
- $\left.\frac{\hbar^{2}\lambda^{2}}{6(4\pi)^{4}}\left[-3\Omega^{2}\left(L_{\Omega}^{2}+3L_{\Omega}+C_{2}\right)+3T^{2}\left(L_{\Omega}J_{1,\Omega}+H_{2,\Omega}+H_{3,\Omega}\right)\right]\right]$
@@
- $\left.\frac{\varphi^{4}}{4!}\frac{3\hbar\lambda^{2}}{2(4\pi)^{2}}\left[L_{\Omega}+J_{2,\Omega}\right]\right\}+\mathcal{F}_{\delta^{2}}\;,$
@@
where we have defined $L_{\Omega}=\ln(\mu^{2}/\Omega^{2})$, with $\mu$ being the renormalization energy scale in the $\overline{\rm MS}$ regularization scheme while $C_{0}$, $C_{1}$ and $C_{2}$ are numerical constants given, respectively, by $C_{0}=-9.8424$, $C_{1}=39.429$ and $C_{2}=3.33288$. The quantities $J_{0,\Omega}$, $J_{1,\Omega}$, $J_{2,\Omega}$, $H_{2,\Omega}$, $H_{3,\Omega}$, $K_{2,\Omega}$ and $K_{3,\Omega}$ are all functions of $\Omega/T$ and their explicit expressions are given in App. A, while the expressions for the terms $\mathcal{F}_{\delta}$ and $\mathcal{F}_{\delta^{2}}$ appearing also in Eq. (6) are provided in App. B. In the expression for the effective potential, Eq. (6), $\varphi$ denotes the constant background scalar field, while $m$ and $\lambda$ represent the renormalized mass and coupling. Note that $m$ and $\lambda$ are implicit functions of the renormalization scale $\mu$. In the next section, we review the approach to improving the effective potential using the renormalization group equation (RGE). Then, in Sec. IV, we show how this approach can be adapted to the OPT and applied to Eq. (6).
In principle, after setting $\delta=1$ one should impose that the OPT effective potential satisfies the RGE
$\left(\mu\frac{\partial}{\partial\mu}+\beta_{\lambda}\frac{\partial}{\partial\lambda}+\gamma_{m}m^{2}\frac{\partial}{\partial m^{2}}+\mu\frac{\partial\eta}{\partial\mu}\frac{\partial}{\partial\eta}+\beta_{\Lambda}\frac{\partial}{\partial\Lambda}-\gamma_{\varphi}\varphi\frac{\partial}{\partial\varphi}\right)\Bigr{|}_{\eta={\overline{\eta}}}V_{\rm OPT}^{\delta^{2}}=0,$ (7)
where $\mu$ is the regularization scale, while $\beta_{\lambda},\,\gamma_{m},\,\beta_{\Lambda}$ and $\gamma_{\varphi}$ are the RGE functions (see the next section). However, due to the PMS equation (5), the term proportional to $(\partial\eta)/(\partial\mu)$ in (7) does not contribute. Now, in principle, the optimized mass $\overline{\eta}$ can be an explicit function of the renormalization scale, as well as of the original parameters, such as $m$, $\lambda$. This suggests that, as far as the RG is concerned, the explicit ${\overline{\eta}}(\mu)$ running is not crucial, so that one should be primarily concerned with the implicit $\mu$-dependence the optimal $\overline{\eta}$ acquires when becoming a function of the original ($\mu$) parameters. Therefore, in order to implement our procedure, we propose that as a second step, one uses the RG to improve the effective potential of the original theory, as we do in the next section.
III Renormalization Group Improvement
In the context of the effective potential of the original theory, the equation for the RGE can be written as [59]
@@ III Renormalization Group Improvement
$\left(\mu\frac{\partial}{\partial\mu}+\beta_{\lambda}\frac{\partial}{\partial\lambda}+\gamma_{m}m^{2}\frac{\partial}{\partial m^{2}}+\beta_{\Lambda}\frac{\partial}{\partial\Lambda}-\gamma_{\varphi}\varphi\frac{\partial}{\partial\varphi}\right)V_{\rm eff}(\varphi)=0,$ (8)
@@
with $Z_{\varphi}$ representing the wavefunction renormalization term. For the $\lambda\phi^{4}$ theory, the functions $\beta_{\lambda}$, $\gamma_{m}$, $\beta_{h}$, and $\gamma_{\varphi}$ are well known and are defined, up to order $\hbar^{3}$, as [60]
@@
$\beta_{0}$ $=$ $\frac{3}{(4\pi)^{2}},\;\;\beta_{1}=-\frac{17}{3(4\pi)^{4}},\;\;\beta_{2}=\frac{1}{(4\pi)^{6}}\left[\frac{145}{8}+12\zeta(3)\right],$
$\gamma_{m0}$ $=$ $\frac{1}{(4\pi)^{2}},\;\;\gamma_{m1}=-\frac{5}{6(4\pi)^{4}},\;\;\gamma_{m2}=\frac{7}{2(4\pi)^{6}},$
@@
$\beta_{\Lambda 0}$ $=$ $\frac{1}{2(4\pi)^{2}},\;\;\beta_{\Lambda 1}=0\;\;\text{, and}\;\;\beta_{\Lambda 2}=\frac{1}{16(4\pi)^{6}}.$ (14)
An effective potential fully satisfying Eq. (8) is invariant under the RGE. In practice, the literature presents some methods that provide ways to obtain an approximate solution to Eq. (8) in a way that mitigates the scale dependence of the effective potential. For our purposes, the method introduced in Ref. [61] and later explored by the authors of Refs. [62, 63, 64, 65, 66] will prove to be the most useful. The improvement of the effective potential through the properties of the renormalization group has also been investigated in the context of finite temperature in Ref. [67] at leading order in the loop expansion. The calculations for the scalar theory $\lambda\phi^{4}$ at $T=0$ were later studied in Refs. [68] and [69] at two- and three-loop orders, respectively. This method is generically known as Renormalization Group Improvement (RGI) of the effective potential.
@@
where the explicit expressions for each term in the above solutions can be found in Refs. [68, 69]. For completeness, we reproduce these solutions in App. C where they are explicitly presented in terms of Eqs. (55), (58), (60)–(69).
@@
In the next section, we will make use of the procedure outlined above for the RGI of the effective potential, but apply it instead to the effective potential derived using the OPT method at finite temperatures.
IV Setting up the Variational Renormalization Group Method
To implement the finite-temperature effective potential derived from the OPT approach with the RGI method, some changes to the original application detailed in the previous section are required for a consistent implementation. We will describe the necessary steps in detail below.
In its original implementation, the RGI method is applied to the effective potential using an $\hbar$-expansion, with the RGE solved order by order in $\hbar$. Therefore, the effective potential can be rewritten as a sum of logarithmic terms whose order matches the $\hbar$ order. In this way, the solutions of Eq. (17) as well as the improved potential are also organized in powers of $\hbar$.
On the other hand, the OPT effective potential is expanded in powers of $\delta$, which essentially tracks the perturbative expansion in $\lambda$ with each order in $\delta$ containing terms of distinct powers of $\hbar$. To reconcile this possible inconsistency, we propose the implementation of an extra step within our prescription. That is, the RGI procedure is applied in powers of $\hbar$ up to the highest found within the OPT effective potential, thereby constraining the coupling order. For example, consider the OPT effective potential at first order in $\delta$, whose highest order in $\hbar$ comes from the double-bubble diagram represented by the second diagram in the first line of Fig. 1 and which is $\mathcal{O}(\hbar^{2})$. Hence, the RGI procedure must be carried out up to order-$\hbar^{2}$; otherwise, if it is applied only up to order-$\hbar$, the contribution from the double-bubble diagram will be lost in the OPT effective potential. In the case of the second-order $\delta^{2}$ potential, the highest $\hbar$-order comes from the basketball diagram, which is given by the last diagram shown in the first line of Fig. 1 and which is $\mathcal{O}(\hbar^{3})$. Hence, at this order of the OPT, the RGI procedure goes up to order-$\hbar^{3}$. This ensures that no information is lost from the original effective potential derived from the OPT, while at the same time, there is a gain of information from the renormalization group. To avoid inconsistencies in the coupling order, i.e., to prevent higher-order terms like $\lambda^{k+1}$ from appearing at an order-$\delta^{k}$, these higher-order coupling terms are discarded. In this way, we can ensure that the effective potential considered in the VRG consistently matches the coupling order of the original OPT effective potential while simultaneously incorporating the RGI method at the appropriate $\hbar$-order. As we will check later through our numerical results, this procedure does ensure a systematic reduction of the scale dependence of the physical quantities that we will be studying. Thus, we find that the solution of the RGE applied to the OPT effective potential is still as given by Eq. (8). Hence, according to the RGI method, we must reinsert the barred parameters given by Eqs.  (18)-(21) into the original OPT effective potential, which must now satisfy the relation
$\!\!V_{\rm eff,OPT}^{\delta^{k}}(\mu,\lambda,m^{2},\phi,\Lambda,\eta,T)\!\!=\!\!V_{\rm eff,OPT}^{\delta^{k}}(\bar{\mu},\bar{\lambda},\bar{m}^{2},\bar{\phi},\bar{\Lambda},\eta,T).$
Equation (IV) also provides us with the boundary condition: as $t\to 0$, we must recover the original effective potential without the RGI improvement. Following the procedure used for the effective potential in the last section using the RGI procedure, except that now we consider the effective potential derived from the OPT, e.g. Eq. (6), substituting in it Eqs. (55), (58), (6069), which, as explained above, is then re-expanded in powers of $\hbar$ up to the highest order contained in the OPT effective potential. In the last step, we finally substitute the solutions given by Eqs. (55), (58), (6069).
Next, fixing the value of $\bar{\mu}$ must be done carefully observing the relevant OPT order. For example, at first order in the OPT, at high temperatures all original logarithms can be rewritten in the form $\ln(\alpha T/\bar{\mu})$, where $\alpha=4\pi/e^{\gamma_{E}}$ is a constant. In the RGI procedure [69], a natural choice for $\bar{\mu}$ is $\bar{\mu}=\alpha T$, so as to control the scale dependence of the scale-dependent log terms. At the next order in the OPT, order-$\delta^{2}$, the same choice can be made as in the previous order. After applying this step-by-step procedure, we obtain the VRG effective potential, e.g., the OPT effective potential improved by RGI. In App. D, we give the explicit expressions for the effective potential that follows from this procedure, with the result at order $\delta$ given by Eq. (70), while the one obtained at order $\delta^{2}$ is given by Eq. (72). Finally, to fix the arbitrary mass parameter $\eta$, we use the PMS criterion, following the standard OPT prescription.
V Results

### v2

@@ I Introduction
In the context of OPT, a new scheme called Renormalization Group Optimized Perturbation Theory (RGOPT) was developed to address the scale dependence problem [41, 42, 43, 44, 45, 46, 47, 48]. This method modifies the original OPT implementation and introduces a new (subtraction) term into the free energy, ensuring invariance under the renormalization group. The traditional OPT mass term, in a scalar field theory, $(1-\delta)\eta^{2}$, is written as $(1-\delta)^{a}\eta^{2}$ where $a$ turns out to be a function of the RG $\beta$ and $\gamma_{m}$ coefficients $b_{0}$ and $\gamma_{0}$. These adaptations have allowed this approximation to achieve great success in investigating the symmetric phase of several models, including QCD, as mentioned earlier. However, so far the method has not yet been used, for example, in the description of phase transitions. The approach that we propose in the present paper differs fundamentally from the RGOPT by preserving the original prescription of the OPT method and improving over the scale dependence displayed by the thermodynamic quantities by applying directly the OPT resummation over the RGI improved effective potential. We termed this alternative approach the Variational Renormalization Group (VRG). We will see that by preserving the original OPT method, augmented with RG invariance properties, we can study the system’s thermodynamic properties in both the symmetric and broken phases in a systematic way.
This work is organized as follows. In Sec. II, we briefly review the application of the OPT method to our fiducial $\lambda\phi^{4}$ scalar field theory. The effective potential at finite temperature up to second order in the OPT is explicitly written down. In Sec. III, we review how the RGI method allows an improvement of the effective potential with respect to scale variations. We also show in that section how the RGI method can be modified to carry out an improvement in the context of the OPT finite-temperature effective potential, ensuring that the improvement remains consistent with the renormalization group. In Sec. IV, we present our main numerical results for the thermodynamics of the quartic scalar field when using the VRG. Finally, in Sec. V, we draw our main conclusions and perspectives for future work. Five appendices are also included where the relevant definitions, expressions and technical details used in our analysis are given.
@@ II The effective potential of the $\lambda\phi^{4}$ model in the OPT at next-to-leading order
$\mathcal{L}$ $=$ $\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi)-\frac{m^{2}}{2}\phi^{2}-\frac{\lambda}{4!}\phi^{4}.$ (1)
The OPT prescription starts by implementing in the Lagrangian density the following replacements [29, 30, 31, 32]
@@
$\mathcal{L}^{\delta}$ $=$ $\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi)-\frac{\Omega^{2}}{2}\phi^{2}+\frac{\delta\eta^{2}}{2}\phi^{2}-\frac{\delta\lambda}{4!}\phi^{4},$
@@
In practice, when applying the OPT, one can use the results obtained by standard perturbation theory up to some order $\lambda^{k}$ in the coupling and then apply the changes given by Eqs. (2) and (3), while re-expanding the result up to the desired order $\delta^{k}$. In the present application, we are mainly concerned with the finite temperature effective potential for the scalar quartic model at order-$\lambda^{2}$. Applying Eqs. (2) and (3) to the perturbative ${\cal O}(\lambda^{2})$ standard result and re-expanding yields the corresponding OPT expression up to order-$\delta^{2}$ (NLO), whose contributions are represented by the Feynman diagrams shown in Fig. 1.
@@
$V_{\rm OPT}=\Delta V_{\rm OPT}^{\delta^{0}}\delta^{0}+\Delta V_{\rm OPT}^{\delta^{1}}\delta^{1}+\Delta V_{\rm OPT}^{\delta^{2}}\delta^{2}+{\cal O}(\delta^{3}),$ (6)
where
$\Delta V_{\rm OPT}^{\delta^{0}}$ $=$ $\frac{1}{2}\left[m^{2}+\eta^{2}\right]\varphi^{2}$ (7)
- $\frac{\Omega^{4}\left(2L_{\Omega}+3\right)}{8\left(4\pi\right)^{2}}-\frac{J_{0,\Omega}T^{4}}{2\left(4\pi\right)^{2}},$
$\Delta V_{\rm OPT}^{\delta^{1}}$ $=$ $-\frac{1}{2}\eta^{2}\varphi^{2}+\frac{\lambda}{4!}\varphi^{4}$ (8)
$+$ $\frac{\hbar\eta^{2}}{2\left(4\pi\right)^{2}}\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]$
- $\frac{\varphi^{2}}{2}\frac{\lambda}{2(4\pi)^{2}}\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]$
$+$ $\frac{\lambda}{8\left(4\pi\right)^{4}}\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]^{2}$
$+$ $\mathcal{F}_{\delta},$
and
$\Delta V_{\rm OPT}^{\delta^{2}}$ $=$ $-\frac{\eta^{4}}{4\left(4\pi\right)^{2}}\left[L_{\Omega}+J_{2,\Omega}\right]-\frac{\lambda}{4\left(4\pi\right)^{4}}\eta^{2}\left[L_{\Omega}+J_{2,\Omega}\right]\left[\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right]$ (9)
- $\frac{\lambda^{2}}{48\left(4\pi\right)^{6}}\left[\left(\left(3L_{\Omega}+4\right)J_{1,\Omega}^{2}+J_{1,\Omega}^{2}J_{2,\Omega}+2K_{2,\Omega}+\frac{4}{3}K_{3,\Omega}\right)3T^{4}\right$.
@@
$+$ $\frac{\varphi^{2}}{2}\left[-\frac{\lambda^{2}\Omega^{2}}{4(4\pi)^{4}}\left((L_{\Omega}+1)^{2}+1+\frac{\pi^{2}}{6}\right)+\frac{\lambda\eta^{2}}{2(4\pi)^{2}}\left(L_{\Omega}+J_{2,\Omega}\right)+\frac{\lambda^{2}}{4(4\pi)^{4}}\left(\left(L_{\Omega}+1\right)\Omega^{2}-J_{1,\Omega}T^{2}\right)\left(L_{\Omega}+J_{2,\Omega}\right)\right$.
@@
- $\left.\frac{\lambda^{2}}{6(4\pi)^{4}}\left[-3\Omega^{2}\left(L_{\Omega}^{2}+3L_{\Omega}+C_{2}\right)+3T^{2}\left(L_{\Omega}J_{1,\Omega}+H_{2,\Omega}+H_{3,\Omega}\right)\right]\right]-\frac{\varphi^{4}}{4!}\frac{3\lambda^{2}}{2(4\pi)^{2}}\left[L_{\Omega}+J_{2,\Omega}\right]+\mathcal{F}_{\delta^{2}}\;,$
@@
where we have defined $L_{\Omega}=\ln(\mu^{2}/\Omega^{2})$, with $\mu$ being the renormalization energy scale in the $\overline{\rm MS}$ regularization scheme while $C_{0}$, $C_{1}$ and $C_{2}$ are numerical constants given, respectively, by $C_{0}=-9.8424$, $C_{1}=39.429$ and $C_{2}=3.33288$. The quantities $J_{0,\Omega}$, $J_{1,\Omega}$, $J_{2,\Omega}$, $H_{2,\Omega}$, $H_{3,\Omega}$, $K_{2,\Omega}$ and $K_{3,\Omega}$ are all functions of $\Omega/T$ and their explicit expressions are given in App. A, while the expressions for the terms $\mathcal{F}_{\delta}$ and $\mathcal{F}_{\delta^{2}}$ appearing also in Eq. (6) are provided in App. B. In the expression for the effective potential, Eq. (6), $\varphi$ denotes the constant background scalar field, while $m$ and $\lambda$ represent the renormalized mass and coupling. Note that $m$ and $\lambda$ are implicit functions of the renormalization scale $\mu$ and they are obtained from the loop expansion of the original theory, prior to the application of the OPT prescription given by (2) and (3).
The next step is the implementation of the RGI procedure to the the original loop expanded $V_{\rm eff}$ (see App. C). Once the relevant RGI $V_{\rm eff}$ is determined one may apply the OPT procedure in order to resum the perturbative series. In the next section we review these steps, starting by determining the RGI $V_{\rm eff}$.
III Renormalization Group Improvement and the OPT Implementation
In this section, we present a brief summary of the method that allows for the mitigation of scale dependence in an effective potential within perturbation theory in $\hbar$. In the context of the effective potential of the massive $\lambda\phi^{4}$ theory at zero temperature $V_{\rm eff}$ which explicitly expression can be found in Ref. [58], the effective potential must satisfied the Renormalizations Group Equation (RGE), that can be written as [59]
@@ III Renormalization Group Improvement and the OPT Implementation
$\left[\mu\frac{\partial}{\partial\mu}+\beta_{\lambda}\frac{\partial}{\partial\lambda}+\gamma_{m}m^{2}\frac{\partial}{\partial m^{2}}+\beta_{\Lambda}\frac{\partial}{\partial\Lambda}-\gamma_{\varphi}\varphi\frac{\partial}{\partial\varphi}\right]V_{\rm eff}(\varphi)=0,$ (10)
@@
with $Z_{\varphi}$ representing the wavefunction renormalization term. Note that to implement the RG procedure, we also introduce the “cosmological constant” $\Lambda$ [60]. In the RGI context, $\Lambda$ is used to ensure the scale invariance of the effective potential [61, 62, 63]. For the $\lambda\phi^{4}$ theory, the functions $\beta_{\lambda}$, $\gamma_{m}$, $\beta_{\Lambda}$, and $\gamma_{\varphi}$ are well known and are defined, up to order $\hbar^{3}$, as [64]
@@
$\beta_{0}$ $=$ $\frac{3}{(4\pi)^{2}},\;\;\beta_{1}=-\frac{17}{3(4\pi)^{4}},\;\;\beta_{2}=\frac{\frac{145}{8}+12\zeta(3)}{(4\pi)^{6}},$
$\gamma_{m_{0}}$ $=$ $\frac{1}{(4\pi)^{2}},\;\;\gamma_{m_{1}}=-\frac{5}{6(4\pi)^{4}},\;\;\gamma_{m_{2}}=\frac{7}{2(4\pi)^{6}},$
@@
$\beta_{\Lambda_{0}}$ $=$ $\frac{1}{2(4\pi)^{2}},\;\;\beta_{\Lambda_{1}}=0\;\;\text{, and}\;\;\beta_{\Lambda_{2}}=\frac{1}{16(4\pi)^{6}}.$ (16)
An effective potential fully satisfying Eq. (10) is invariant under the RGE. In practice, the literature presents some methods that provide ways to obtain an approximate solution to Eq. (10) in a way that mitigates the scale dependence of the effective potential. For our purposes, the method introduced in Ref. [61] and later explored by the authors of Refs. [62, 63, 65, 66, 67] will prove to be the most useful. The improvement of the effective potential through the properties of the renormalization group has also been investigated in the context of finite temperature in Ref. [68] at leading order in the loop expansion. The calculations for the scalar theory $\lambda\phi^{4}$ at $T=0$ were later studied in Refs. [69] and [58] at two- and three-loop orders, respectively.
@@
where the explicit expressions for each term in the above solutions can be found in Refs. [69, 58]. For completeness, we reproduce these solutions in App. D where they are explicitly presented in terms of Eqs. (59), (62), (64)–(73).
III.1 Obtaining the effective potential in the RGI method
@@
In the context of the finite-temperature effective potential, the same procedure can be applied. New terms arise due to the finite-temperature procedure; these terms also contain the parameters of the theory, which must undergo the appropriate modifications due to the RGI, for further details, see Ref. [68].
@@
Next, we describe procedure outlined above for the RGI of the effective potential in the context of the OPT method at finite temperatures.
III.2 Obtaining the effective potential in the VRG method
We now present a consistent implementation of the VRG method for the finite-temperature effective potential derived within the OPT. The strategy is to apply OPT directly to the RGI finite-temperature effective potential, thereby combining variational resummation with a systematic reduction of renormalization-scale dependence.
The construction proceeds as follows. First, we consider the loop expanded finite-temperature effective potential. The RGI procedure described above is then applied to this effective potential. At this stage, the RG equations are solved without fixing the optimized scale $\bar{\mu}$. This procedure leads to the RGI effective potential as described above.
Next, the OPT deformation, defined by Eqs. (2) and (3), is implemented in the RGI effective potential. When expanding the effective potential in powers of the bookkeeping parameter $\delta$, the quantity $\xi=1-\beta_{0}\lambda t$ is left unexpanded. This prescription ensures that the $\delta$ expansion does not spoil the logarithmic resummation achieved by the RGI method, in close analogy with the treatment of the $\hbar$ expansion in Refs. [69, 58].
The optimum scale $\bar{\mu}$ is next fixed according to the OPT order under consideration. For instance, at high temperatures and at first order in $\delta$, all logarithmic terms can be cast in the form $\ln(\alpha T/\bar{\mu})$, where $\alpha=4\pi/e^{\gamma_{E}}$. This allows us to choose for example $\bar{\mu}=\alpha T$, and adopt the same prescription at order $\delta^{2}$. The resulting VRG effective potentials at orders $\delta$ and $\delta^{2}$ are given in Eqs. (74) and (78), respectively, with explicit expressions collected in App. E. The final step is to obtain the arbitrary mass parameter $\eta$, which is then fixed by applying the PMS condition Eq. (5) to the VRG effective potential.
IV Results